what is work energy thereom...................................?

Grade:12

17 Answers

SAI SARDAR
7 years ago
dear tushan,
                              The work done by the resultant force acting on a body is equal to the change in its kinetic energy.
tushan
7 years ago
dear tushan, The work done by the resultant force acting on a body is equal to the change in its kinetic energy.
raju
7 years ago
Hi ! Tu
Shreyas Patil
7 years ago
Relation bewteen KE and W: The work done on an object by a net force equals the change in kinetic energy of the object: W = KEf - KEi. This relationship is called the work-energy theorem.
rahul kushwaha
7 years ago
For any net force acting on a particle moving along any curvilinear path, it can be demonstrated that its work equals the change in the kinetic energy of the particle by a simple derivation analogous to the equation above.calls by the name of work energy theorem.....
Ankur Saha
7 years ago
To change the kinetic energy of a particle we have to apply a force on it and the force must do work on it. This relationship between KE and WORK is called "work -energy theorem "...[W=KEf-KEi]
Pranav Dabhade
7 years ago
work done by all forces =change in K.E it can also written as,[W=KEf-KEi] .That means kinetic energy lost is converted inti work &vice a versa.,
Anshul
7 years ago
kinetic energy of a particle we have to apply a force on it and the force must do work on it. This is called "work -energy theorem Work = K.Ei - K.Ef
Pavan Kumar Reddy
7 years ago
The amount of work done by the net force acting on a particle is equal change in it’s kinetic energy.
YASWANTH KISHOR
7 years ago
work energy therom is nothing but the total work done by all forces is equal to the change in its kinetic energy …. W=1/2 {kx22-kx12}
poooola dinesh
7 years ago
Total workdone by the Net Force on the body is equal to change in it’s kinetic energy
Wtotal=1/2{kx22-kx12}
Sushmit Trivedi
7 years ago
The work W done by the net force on a particle equals the change in the particle`s kinetic energy KE: W = Δ K E = 1 2 m v f 2 − 1 2 m v i 2 . The work-energy theorem can be derived from Newton`s second law. Work transfers energy from one place to another or one form to another. thank you
jyoti bhatia
7 years ago
The work-energy theorem is a generalized description of motion which states that work done by the sum of all forces acting on an object is equal to change in that object’s kinetic energy. This principle of work and its relationship to kinetic energy is a core mechanical physics concept.
The work-energy theorem describes direct relationship between work and energy, its formula shows that net amount of work done by forces acting on a particle causes a change in that particle’s kinetic energy.
Work is an action that is performed on an object or system and that transfers energy from one location or from one form to another. Basic formula is work = force x distance
Kinetic energy is the energy of motion that an object or system has at a given point in time. Basic mathematical formula is kinetic energy = ½* mass * velocity².
Since work = energy, formulas for each can be substituted to create following work-energy equation: force * distance = ½ * mass * velocity². This equation is transformed into a set of other equations called ‘kinematics equations’.
The work-energy theorem can be derived from Newton’s second law of motion, which concerns inversely proportional relationship between acceleration and mass and directly proportional relationship between net force and acceleration.
Ajay Panthagani
7 years ago
work energy theorem describes relation between the work done by a force on a point object and the change in its kinetic energy.
Wtot=K2-K1= ∆Kp { margin-bottom: 0.25cm; direction: ltr; line-height: 120%; text-align: left; }

 

Alish
7 years ago
It states that,”the work done by application of a force on a body equal to the change in kinetic energy of a body.”
 
Ragini
7 years ago
Acc to this theorem the work done by constant force is equal to the change in KE or increase. in KE.......
Rajdeep
7 years ago
I